3x-6y=10 and 9x+15y=-14
1
Expert's answer
2021-08-02T13:54:13-0400

We should determine such values (x,y) that satisfy both equations:

"\begin{cases}
3x-6y=10, \\ 9x+15y=-14
\end{cases}"

Let us multiply the first equation by 3 and subtract the second equation from the first:

"\begin{cases}
9x-18y=30, \\ 9x+15y=-14
\end{cases}" , "-33y = 44 \Rightarrow y = -\dfrac43\,."

Next, we substitute the value of y into the first equation and obtain the value of x:

"x = \dfrac19(30 + 18y) = \dfrac19(30- 24) = \dfrac23."

Therefore, "(x,y) = \left( \dfrac23, -\dfrac43\right)."